OpenStudy (anonymous):

In professional baseball, a commonly used statistic is winning percentage. To calculate a team's winning percentage, the number of wins is divided by the total games played. Although called a percentage, the winning percentage is usually displayed in decimal form, rounded to three decimal places, like this, for example: .244. The Tacoma Tigers, a new MLB team, were in last place in their division with a record of 29 wins and 43 losses. In other words, their current winning percentage was 29/29+43=29/72=403

OpenStudy (anonymous):

We are going to find out how many games in a row the Tigers have to win to increase their winning percentage to .500. To set up an equation, consider this: The number of games won in a row would have to be added to both the number of wins and the total number of games, and the result would have to be .500.

OpenStudy (anonymous):

I got 29+x/72+x= .5 AND now it is asking for me to solve to determine how many games the Tigers need to win in a row

OpenStudy (anonymous):

U need to fine x?

OpenStudy (anonymous):

Find*

OpenStudy (anonymous):

Well it is asking to solve so I guess haha.... :/

OpenStudy (anonymous):

x is 14

OpenStudy (anonymous):

Okay that is what I got. So the next question is: Choose the equation needed to find the solution to the following question: If the Tigers won 2 out of every 3 (or 2/3) of their remaining games, how many games would it take them to reach a winning percentage of .500?

OpenStudy (anonymous):

21

OpenStudy (anonymous):

They only have to win 14 games. Not 28

OpenStudy (anonymous):

Let y be the remaining number of games. So, 2/3 * y = 14. They, y=21.

OpenStudy (anonymous):

Oh Im sorry I didn't give you any equations to choose from. A. 29+x/72+x= .5 B. 29+2/3x/72+2/3x= .5 C. 29+x/72+2/3x= .5 D.29+2/3x/72+x= .5

OpenStudy (anonymous):

I think it's B

OpenStudy (anonymous):

I actually just looked at the next question and it basically gives away that the answer is D..... The next question is: The equation would be easier to solve if the left-hand side of the equation were simplified from a complex fraction to a standard fraction. Simplify the complex fraction 29+2/3x/72+x

OpenStudy (anonymous):



OpenStudy (anonymous):

Wait I am confused.... so is that simplifying the complex fraction of 29+2/3x/72+x?

OpenStudy (anonymous):

I think... I just checked the link that other guy replied and it has the same answer too.

OpenStudy (anonymous):

Okay well thank you for you help!!! (:

OpenStudy (anonymous):

You're welcome.